OpenStudy (anonymous):

42 is _____% of 140 12 is 15% of (what number)______? 76 is 40% of (what number)______?

OpenStudy (anonymous):

Same principle of= times % = x/100

OpenStudy (anonymous):

x/100 times 140 = 42

OpenStudy (anonymous):

= 30%

OpenStudy (anonymous):

and 15/100 times x = 12 =80

OpenStudy (anonymous):

and 76 is 40% of (what number)______? 40/100 times x = 76 See if you can work this one out.

OpenStudy (anonymous):

im still confused.